A is eleven times as old as B. in a certain number of years, A will be five times as old as B, and in five years after that she will be three times as old as B. how old are they now.: A is eleven times as old as B. in a certain number of years, A will be five times as old as B, and in five years after that she will be three times as old as B. how old are they now.
Answer by stanbon(18792)

A is eleven times as old as B.
A = 11B
-------------
in a certain number of years, A will be five times as old as B
A+x = 5(B+x)
-------------
and in five years after that she will be three times as old as B.
A+x+5 = 3(B+x+5)
-----------------
how old are they now?
-------
Rearrange the three equations with variables A, B, and x
A - 11B + 0x = 0
A - 5B -4x = 0
A -3B - 2x = 10
---------------
Solve the systems of three equations by any method to get:
A = 22
B = 2
x = 3
-----------
A is 22 yrs. old
B is 2 yrs. old
======================
Chers,
Stan H.